### v2

# Document Reconstruction Unlocks Scalable Long-Context RLVR

Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) has become a prominent paradigm to enhance the capabilities (i.e. long-context) of Large Language Models (LLMs). However, it often relies on gold-standard answers or explicit evaluation rubrics provided by powerful teacher models or human experts, which are costly and time-consuming. In this work, we investigate unsupervised approaches to enhance the long-context capabilities of LLMs, eliminating the need for heavy human annotations or teacher models’ supervision. Specifically, we first replace a few paragraphs with special placeholders in a long document. LLMs are trained through reinforcement learning to reconstruct the document by correctly identifying and sequencing missing paragraphs from a set of candidate options. This training paradigm enables the model to capture global narrative coherence, significantly boosting long-context performance. We validate the effectiveness of our method on two widely used benchmarks, RULER and LongBench v2. While acquiring noticeable gains on RULER, it can also achieve a reasonable improvement on LongBench v2 without any manually curated long-context QA data. Furthermore, we conduct extensive ablation studies to analyze the impact of reward design, data curation strategies, training schemes, and data scaling effects on model performance. We publicly release our code, data, and models [^1] https://github.com/XYaoooo/reconstruction_long.
1 Introduction
Figure 1: We report the average score of RULER and overall score of LongBench v2 for Qwen2.5-7B-Instruct-1M and LLaMA-3.1-8B-Instruct.
Reinforcement Learning with Verifiable Rewards (RLVR) has recently achieved the state-of-the-art in Large Language Model (LLM) reasoning (Kumar et al., 2025; Yu et al., 2025; Yeo et al., 2025; Zeng et al., 2025a; Liu et al., 2025; Anthropic, PBC, 2025). Using ground-truth feedback to guide generation, RLVR enables models like DeepSeek-R1 (Guo et al., 2025) to navigate complex multi-step problem-solving paths in domains such as mathematics and programming with unprecedented precision (Yang et al., 2025a). However, as LLMs evolve into agents that must interact with expansive real-world datasets, the challenge shifts from local reasoning to global context processing (Team et al., 2025b; c; Prabhakar et al., 2025). Here, a significant gap remains: models that excel at step-by-step logic reasoning often struggle to maintain coherence when retrieving and synthesizing information across tens of thousands of tokens (Wu et al., 2025b; Zhuang et al., 2025; Wu et al., 2025c; Bai et al., 2025b). This suggests that reasoning gains do not necessarily scale with context length, leaving a gap between reasoning capacity and long-context understanding.
The pursuit of extending the context window has thus become a central objective in developing frontier LLMs (Zhu et al., 2024; An et al., 2024; Peng et al., 2024; Gao et al., 2025; Lu et al., 2025; Yang et al., 2025b; Wan et al., 2025; Xu et al., 2026). From analyzing massive code bases to distilling insights from an entire document, the ability to process and reason over long sequences is essential for practical applications. However, as the context length increases, the difficulty of maintaining global coherence and performing precise retrieval increases exponentially. Models frequently suffer from the well-documented “lost in the middle” phenomenon (Liu et al., 2024), where information in the center of a long prompt is ignored, or they struggle to maintain the logical consistency of a narrative over tokens (Hsieh et al., 2024b; a; Du et al., 2025). Although RLVR offers a powerful framework for LLMs to refine their long-range dependency handling, its application is currently restricted by a heavy reliance on external supervision.
Recent RLVR approaches (Wang et al., 2025; Chen et al., 2026) to enhance the long-context capability of LLMs are constrained by the availability of gold-standard answers or evaluation rubrics, which are typically provided by expensive human experts and frontier teacher models (often close-source) (Chen et al., 2025a; Zhang et al., 2025; Huang et al., 2025). This dependency creates a significant scalability bottleneck: while the need for long-context understanding is universal, question-answering pairs from human annotations required for training are prohibitively expensive and difficult to generate at the scale needed for reinforcement learning. Furthermore, relying on teacher models can introduce biases or limit the potential of student models to the capabilities of the supervisor (Kim et al., 2025; Cheng and Amiri, 2025). To unlock the next level of long-context ability of LLMs, we explore a training mechanism that can derive objective, verifiable rewards directly from the data itself in an unsupervised manner.
In this work, we introduce a fully unsupervised RLVR framework that bypasses the need for costly human annotations or rubrics from teacher models, enabling a more scalable approach to long-context training. We hypothesize that long documents possess an inherent structure, specifically their narrative flow and logical sequence, which contains valuable internal signals. The signals can act as a natural and verifiable reward for training model through RLVR. We formalize this through a document reconstruction task. We first mask a few random paragraphs within a long document and require the LLM to correctly identify and sequence these missing segments from a shuffled pool of candidates. Since the original documents provide the ground truth, the resulting reward enjoys the desirable property of being both verifiable and fully automated. This reconstruction training mechanism encourages LLMs to move beyond surface-level pattern matching and instead develop a deeper, more structural understanding of the global context (Yang et al., 2019).
We evaluate our method on two of the most rigorous benchmarks in the field: RULER (Hsieh et al., 2024a) and LongBench v2 (Bai et al., 2025a). Our empirical results demonstrate that this unsupervised paradigm yields substantial gains on RULER and moderate improvement performance on LongBench v2, as shown in Figure 1. These findings suggest that the underlying long-range structural integrity of documents provides a valuable training paradigm, offering a scalable path toward more capable long-context LLMs. Our contributions are summarized as follows:
  • We propose an unsupervised RLVR framework based on document reconstruction to improve the long-context capabilities of LLMs, eliminating the need for human annotations or teacher model supervision.
  • We formulate document reconstruction as a sequential selection problem with verifiable rewards, encouraging models to learn global narrative coherence and long-range structural dependencies.
  • We validate our approach through extensive experiments on RULER and LongBench v2, showing that it provides a scalable and effective alternative to supervised long-context training.
  • We perform comprehensive ablation studies to examine the effects of reward formulation, data curation strategies, and training configurations, offering deeper insights into the characteristics of our method.
2 Background
Group Relative Policy Optimization (GRPO).
Group Relative Policy Optimization (GRPO) (Shao et al., 2024; Guo et al., 2025) is a policy-gradient method that removes the need for an explicit value function by estimating advantages through relative comparisons within a group of sampled responses. Given a query $q$, GRPO samples a group of trajectories $\{\tau^{(i)}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)$ and assigns each trajectory a scalar reward $R_{\tau^{(i)}}$. The advantage for each trajectory is computed by normalizing its reward against the group:
$A(\tau^{(i)})=\frac{R_{\tau^{(i)}}-\operatorname{mean}\!\left(\{R_{\tau^{(j)}}\}_{j=1}^{G}\right)}{\operatorname{std}\!\left(\{R_{\tau^{(j)}}\}_{j=1}^{G}\right)}$
GRPO then optimizes a PPO-style clipped surrogate objective using these group-relative advantages, enabling stable policy updates without learning a separate value function.
Using the normalized group-relative advantages, GRPO performs policy updates via a PPO-style clipped surrogate objective:
$\mathcal{L}_{\text{GRPO}}(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta_{\text{old}}}}\Big[\min\!\big(r_{\theta}(\tau)A(\tau),\operatorname{clip}(r_{\theta}(\tau),1-\epsilon,1+\epsilon)A(\tau)\big)\Big]$
where
$r_{\theta}(\tau)=\frac{\pi_{\theta}(\tau\mid q)}{\pi_{\theta_{\text{old}}}(\tau\mid q)}.$
This clipped objective constrains policy updates while leveraging group-normalized rewards as a low-variance advantage estimator, eliminating the need for a learned critic.
3 Method
Figure 2: Overview of the document reconstruction framework. Given a long document, we corrupt it by selecting some paragraphs and shuffle them as options. We train LLMs via RLVR to reconstruct the document by generating the option sequence by order.
3.1 Task Formulation: Document Reconstruction
Our goal is to derive a verifiable reward from raw text without labels from human annotators. Given a long document consisting of $n$ paragraphs, denoted as $D=\{p_{1},p_{2},\ldots,p_{n}\}$, we first select a subset of paragraphs to mask. The original document is then transformed into a corrupted context, where the masked paragraphs are replaced by placeholders marked with identifiers, denoted as <CHUNK_i>MISSING</CHUNK_i>. $i$ means the $i$-th masked paragraph of the document.
The model is presented with the context and a set of shuffled candidates labeled with options. LLMs are asked to reconstruct the original text by first reasoning and then generating a list of these options in the correct order. For example, if four paragraphs were selected, the model’s output would be a formatted list such as $\{B,A,D,C\}$. This formulation transforms the long context understanding problem into a sequential decision making task, where the model must utilize global narrative flow and logical consistency to determine the correct placement of each segment. The overview of our method can be found in Figure 2.
3.2 Reward Design
Unlike open-ended generation, our reconstruction task provides an objective and verifiable answer. Since the original document provides ground truth ordering, we define a verifiable reward function that evaluates the model output against ground truth as follows:
$\mathrm{Reward}(o,g)=\begin{cases}1,&\text{if }o=g,\\[4.0pt] \dfrac{1}{K}\sum_{i=1}^{K}\mathbb{I}\!\left[o_{i}=g_{i}\right],&\text{if }\mathrm{ValidPermutation}(o)\land o\neq g,\\[6.0pt] 0,&\text{otherwise}.\end{cases}$ (1)
In this formulation, $K$ represents the total number of masked segments, while $o_{i}$ and $g_{i}$ denote the predicted and ground-truth options for the $i$-th placeholder, respectively. The function $\mathbb{I}[\cdot]$ is an indicator function that yields $1$ for a correct match and 0 otherwise. A critical component of this reward structure is the $\mathrm{ValidPermutation}(o)$ condition. We define a predicted output as a valid permutation if and only if the set of options provided in the model’s response is identical to the set of ground-truth options. This constraint requires the model to correctly identify and utilize the complete pool of candidates, without omitting or duplicating any options (Wu et al., 2025a; Lu et al., 2026).
This piecewise reward structure balances global accuracy with fine-grained feedback. A full reward of 1 is assigned when the reconstruction is exact ($o=g$). For outputs that are not perfectly ordered but remain structurally valid, defined as sequences whose predicted option set exactly matches the ground truth, we assign a partial reward proportional to the fraction of correctly placed segments. This design encourages the model to progressively refine its global structure even when full sequence accuracy is not achieved. In contrast, any output that violates the required format or constitutes an invalid permutation receives a reward of 0. The entire training signal is fully automated and does not depend on human annotations or external teacher models.
3.3 Curriculum through Complexity Scaling
An appealing advantage of our framework is the ability to precisely calibrate the difficulty of training samples by adjusting the complexity of the reconstruction task. Our intuition is that larger $K$ leads to more challenging samples. As $K$ increases, the search space expands exponentially because the number of possible permutations for the candidate set is $K!$. By treating $K$ as a tunable hyperparameter, we can implement a curriculum (Bengio et al., 2009) that allows the model to first master local coherence with fewer options before tackling the complex global structural dependencies required for extremely long contexts. This controllable difficulty (Zeng et al., 2025b; Wang et al., 2026a) ensures that the model can progressively build its long-context understanding capabilities.
4 Experimental Setup
Data Curation
We source our training documents from the corpus provided by Chen et al. (2025a), which covers three diverse domains: books, arXiv papers, and code. We curate a subset consisting of the 8,000 longest documents from the book domain, along with 3,000 documents each from arXiv and code (total 14,000). We apply varying levels of difficulty to these documents by setting $K\in\{2,4,6,8\}$ with their corresponding number ratio being $3:3:3:5$, which we find useful in our preliminary experiment. The average token number of samples in the training set is 49,000. This multi-scale approach to the number of masked segments ensures that the training set provides a broad spectrum of structural challenges. In addition, we also curate 500 samples for validation set to better observe training dynamics.
Training.
We adopt the curriculum schedule (Bengio et al., 2009) to progressively train models by increasing $K$ from 2 to 8. Our implementation is built on the Verl framework (Sheng et al., 2024). To optimize the training process, we employ Group Relative Policy Optimization (GRPO) (Shao et al., 2024). For RL training, we use the AdamW optimizer with a constant learning rate of 1e-6 and a 5-step linear warmup. During rollouts, we use a prompt batch size of 128 and sample 8 responses per prompt, with a maximum context length of 64K and a response length of 4096. Our reconstruction prompt can be found below.
[…]
Evaluation.
Benchmarks. We evaluate all models on two challenging long-context QA benchmarks: (1) RULER (Hsieh et al., 2024a): A synthetic benchmark testing multi-hop reasoning over arbitrary context length. Specifically, we evaluate on four tasks of it (Variable Tracking, Frequent Words Extraction, Common Words Extraction, Question Answering). (2) LongBench v2 (Bai et al., 2025a): A realistic multi-choice QA benchmark on documents up to 128K tokens. We evaluate the lengths of 32K, 64K, and 128K. In our analysis, we primarily focus on RULER, as it provides a more comprehensive and diverse task results.
Models and Baselines.
We select LLaMA-3.1-8B-Instruct and Qwen2.5-7B-Instruct-1M as our backbone models, which also serve as our baseline. A discussion about continual pretraining on long documents is also presented in Appendix A.1.
5 Results and Analysis
5.1 Main Results
(a) Qwen2.5-7B-Instruct-1M
(b) LLaMA-3.1-8B-Instruct
Figure 3: Performance comparison across different context lengths and models.
We summarize the main results across different context lengths and backbone models in Figure 3. In addition, we also report the average on RULER and the overall score in Figure 1.
On RULER, our method produces substantial improvements, with consistent gains as the context length increases from 32K to 128K. This trend indicates that our unsupervised reconstruction training effectively enhances the model’s ability to maintain global coherence and retrieve relevant information in long contexts. On LongBench v2, we also observe moderate improvements across most context lengths. And these gains are achieved without using any manually curated long-context QA data, highlighting the effectiveness of reconstruction training even for question answering. The improvements remain consistent across different backbone architectures, including Qwen2.5-7B-Instruct-1M and LLaMA-3.1-8B-Instruct, demonstrating that our method is not tied to a specific model family.
Taken together, these results show that unsupervised document reconstruction via RLVR provides a scalable and effective mechanism for improving long-context capabilities. The method delivers strong gains on synthetic reasoning benchmarks and meaningful improvements on realistic QA tasks, all while eliminating the need for human annotations or teacher-model supervision. We also record the performance on our curated validation set during the training process in Appendix A.2. Detailed scores on RULER can be found in Appendix A.3.
5.2 Dense vs. Sparse Reward
Figure 4: Average scores of RULER. We compare performance of dense and sparse rewards.
In our main experiments, we employ a dense reward that provides partial credit for partially correct reconstructions. To better understand the role of reward shaping in document reconstruction, we compare this design with a sparse reward formulation, defined as
$\mathrm{Reward}(o,g)=\begin{cases}1,&\text{if }o=g,\\[4.0pt] 0,&\text{otherwise}.\end{cases}$ (2)
This sparse reward assigns a non-zero signal only when the predicted ordering exactly matches the ground truth, providing a stricter but less informative supervision signal. As shown in Figure 4, sparse rewards obtain performance similar to dense rewards on LLaMA-3.1-8B-Instruct. However, it causes significant performance degradation on Qwen2.5-7B-Instruct-1M. We hypothesize that sparse rewards are more likely to cause training instability due to the sparsity of positive rewards in the training process.
5.3 Robustness to Option Mixture Ratios
Figure 5: Average scores of RULER. We compare the performance of different option length mixture ratios.
In our main experiments, we adopt the option mixture ratio for $K=2,4,6,$ and $8$ as $3:3:3:5$, which assigns only a moderate portion of training samples to small values of $K$ (e.g., KK=2). To further validate robustness, we conduct an ablation study by shifting more training samples toward larger $K$, using the ratio $1:2:2:2$. In Figure 5, empirical results demonstrate that our method maintains high performance across varying option length distributions. For the Qwen2.5-7B-Instruct-1M model, the $1:2:2:2$ ratio yielded the highest score, surpassing baseline. In the case of LLaMA-3.1-8B-Instruct model, the $3:3:3:5$ ratio proves most effective. These findings suggest that while specific ratios can offer marginal gains on different model architecture, the overall framework remains robust. The consistency in performance across different difficulty blends confirms that the method does not rely on a brittle or overly specific data composition to succeed.
5.4 Longer Documents Bring Consistent Improvement
Figure 6: We report the average score of RULER
During data curation of main experiment, we retain the longest 8,000, 3,000, and 3,000 documents from the book, arXiv, and code domains, respectively. In this experiment, we compare this document selection strategy against two alternative counterparts: shortest and random. The shortest counterpart selects the shortest documents while preserving the same domain ratios and total number of documents. The random counterpart samples documents randomly, keeping all other factors identical.
As shown in Figure 6, although short-document training brings modest gains for LLaMA-3.1-8B-Instruct, it does not exceed the baseline performance of Qwen2.5-7B-Instruct-1M. Moreover, random document sampling leads to consistent but limited improvements on Qwen2.5-7B-Instruct-1M. In contrast, training on long documents leads to more overall improvements, suggesting the importance of longer contexts for the reconstruction task. In summary, these observations indicate that document length plays a critical role in enabling effective long-context understanding through reconstruction training.
5.5 More Documents, Better Performance
Scaling training data is an important factor in understanding the effectiveness of reconstruction learning. In this part, we study the effect of training data scale on model performance by increasing the number of reconstruction training samples from 0 to 14,000. This setting allows us to examine how model performance evolves when we increase the number of reconstruction samples. For each data scale, we train models under identical optimization settings and evaluate them on RULER. When the data size is set to 0, the model corresponds to the original backbone without reconstruction training, serving as a baseline. We record the performance in Figure 7, model performance generally improves as the number of reconstruction training samples increases.
Figure 7: Average scores of RULER when we scale data from 0 to 14K.
Although performance may fluctuate slightly at smaller data scales, especially when the training set is limited, the overall trend is clearly positive as more data is introduced. In particular, performance consistently increases when the data scale exceeds 4,000 samples, indicating that sufficient reconstruction data is crucial for effectively enhancing long-context understanding. These results suggest that reconstruction-based training scales well with data size and that increasing training data is an effective and stable way to improve model performance. Notably, we do not observe a clear performance plateau within the evaluated data range, suggesting that the model may continue to benefit from additional reconstruction training data. Significant performance gains can be achieved on Qwen2.5-7B-Instruct-1M by scaling data to 3,0000, which can be found in Appendix A.4.
5.6 Training Strategy: Shuffle or Not
Model Qwen LLaMA
Base 68.86 58.76
Shuffle
   True 69.82 65.00
   False 72.20 67.12
Table 1: Effect of training strategy during reconstruction training on RULER.
As mentioned in Section 3, we curate training data by progressively increasing the option length $K$ from 2 to 8, following a curriculum-style training strategy. Our intuition is that reconstruction tasks with larger $K$ are substantially more challenging, and gradually exposing the model to harder samples can facilitate more stable optimization.
To validate this design choice, we compare the curriculum strategy with a shuffled training variant. As shown in Table 1, the curriculum-based strategy (shuffle is false) consistently outperforms shuffled training for both Qwen2.5-7B-Instruct-1M and LLaMA-3.1-8B-Instruct backbones. These results demonstrate that a curriculum of option length is an important component of our training framework.
5.7 Analysis of Option Length
Model Qwen LLaMA
Base 68.86 58.76
$K=2$ 64.79 63.96
$K=4$ 65.83 64.27
$K=6$ 64.69 64.19
$K=8$ 64.83 63.53
Table 2: Average performance of different $K$ values on RULER.
In this part, we study the influence of sample difficulty on model performance. Intuitively, a sample with more candidate options is more challenging, as the model must search over a larger permutation space and rely more heavily on global contextual cues. Specifically, we randomly down-sample 5,000 documents from the training corpus and create four distinct training sets, where each set is constructed using a fixed option length $K\in\{2,4,6,8\}$, without mixing different values of $K$ within the same set. We train an independent model for each curated training set, where each model is associated with a specific $K$.
As shown in Table 2, we observe that model performance is relatively stable across different option lengths $K$. While increasing $K$ substantially enlarges the permutation space and task difficulty, it does not lead to a significant degradation in downstream RULER performance for either backbone. This suggests that the reconstruction-based training objective encourages robust global structure understanding rather than overfitting to a specific difficulty level. Interestingly, intermediate option lengths ($K=4$ and $K=6$) slightly outperform both smaller and larger values of $K$, indicating a potential trade-off between task difficulty and learning efficiency. In addition, training of single $K$ benefits weaker LLaMA-3.1-8B-Instruct, while degrades performance of stronger Qwen2.5-7B-Instruct-1M, suggesting different sensitivity to task specialization across backbones. Qwen2.5-7B-Instruct-1M may be more likely to overfit to a single pattern of reconstruction. It also highlights the importance our data curation strategy in main experiment, which mixes training samples from various $K$.
6 Related Work
Reinforcement Learning with Verifiable Rewards
Reinforcement Learning with Verifiable Rewards provides an objective and scalable framework for improving the reasoning capabilities of large language models by supervising them with ground-truth answers (Lambert et al., 2025; Guo et al., 2025). This paradigm has proven particularly effective in domains with unambiguous correctness criteria, including mathematics and programming. Previous work shows that RLVR can elevate models to expert-level reasoning performance by encouraging the discovery of correct internal reasoning trajectories through outcome-only rewards (OpenAI et al., 2024; Kim et al., 2025; Huang and Yang, 2025; Mo et al., 2025; Wang et al., 2026b). Most existing RLVR studies focus on self-contained reasoning tasks, where the central challenge is to recover a valid trajectory (Yue et al., 2025). In these settings, emergent behaviors, such as self-reflection, have been identified as important contributors to performance gains (Gandhi et al., 2025). Recent reasoning-oriented models, including OpenAI’s o1 series and DeepSeek, have popularized RLVR-based training pipelines and reinforcement learning algorithms such as GRPO.
Long-Context Training of LLMs
Training large language models for long-context reasoning poses challenges that differ fundamentally from those addressed by standard reinforcement learning with verifiable rewards. The outcome reward provides limited guidance in long-context settings, where success depends on identifying and grounding relevant evidence from extensive context (Wan et al., 2025). Most existing approaches to long-context training rely on supervised fine-tuning with synthetic data rather than reinforcement learning (Li et al., 2024; Yen et al., 2025; Chen et al., 2025b). Prior work pads questions with unrelated passages, shuffles document order, or fills contexts with irrelevant text to artificially increase sequence length (Trivedi et al., 2023). However, synthetic-data construction often introduces superficial complexity and fails to produce high-quality and challenging examples, limiting the effectiveness of the resulting long-context capabilities. Despite these limitations, improving long-context reasoning has become increasingly important due to the rapid emergence of agent applications (Zhao et al., 2024; Team et al., 2025a; b; Prabhakar et al., 2025; Gandhi et al., 2026). Thus, developing scalable and principled training methods that move beyond synthetic context expansion and provide effective supervision for long-context training is becoming urgent for advancing large language models.
7 Conclusion
In this work, we present an unsupervised reinforcement learning framework for improving the long-context capabilities of large language models through document reconstruction. By formulating reconstruction as a sequential decision-making problem with verifiable rewards derived directly from raw documents, our approach eliminates the need for manually curated long-context QA data or teacher-model supervision. This enables a scalable and principled alternative to existing long-context training paradigms.
Extensive experiments on RULER and LongBench v2 demonstrate that reconstruction-based RLVR effectively enhances long-context performance across multiple backbone models and context lengths. Beyond empirical gains, our findings highlight document structure itself as a valuable and underexplored supervision signal for long-context learning. We hope that this work motivates further research into unsupervised and self-supervised training objectives that leverage intrinsic structure in raw data and contributes to the development of more capable and scalable long-context language models.
8 Limitations
Despite its effectiveness, our approach has several limitations. Our method requires access to sufficiently long and well-structured documents, which may limit its applicability in domains where long-form data is scarce or noisy. In addition, we observe that the benefits of reconstruction training vary across backbone models. And models may have different requirements about document quality and length. Finally, while our method scales well within the evaluated data range, its behavior at substantially larger scales and with different model sizes remains an open area for future work.
Appendix A Appendix
A.1 On continual Pretraining
We further explore continual pretraining by training the models for one epoch on our curated long-document corpus. However, this approach results in significantly worse performance compared to the original models. We attribute this degradation to two main factors. First, the quality of our collected long documents may not exceed that of the proprietary data used during the original pretraining of the models. Second, continual pretraining on instruction-tuned models may disrupt their instruction-following capabilities. As a result, continual pretraining of LLaMA-3.1-8B-Instruct and Qwen2.5-7B-Instruct-1M does not yield performance improvements.
A.2 On Validation
During training, we monitor model performance on a held-out validation set to assess optimization stability and learning dynamics in reconstruction training. Specifically, we track three metrics: (1) the success rate of answer extraction (i.e., producing a valid permutation), (2) the dense reward, and (3) the sparse reward for Qwen2.5-7B-Instruct-1M. As shown in Figure 8, all three metrics improve smoothly over training steps, indicating stable optimization without severe oscillation or collapse. Notably, the dense reward increases earlier and more steadily than the sparse reward. Overall, these validation trends suggest that the proposed reconstruction-based RLVR framework provides a stable and effective training.
Figure 8: We report metrics on validation set during training process.
A.3 Ruler Score Details
As shown in Table 3, we report the evaluation scores for each subtask for RULER.
Model Task 32k 64k 128k
base ours base ours base ours
Qwen vt 58.32 67.36 46.80 61.56 54.84 56.28
cwe 87.92 90.66 82.46 84.38 69.78 73.42
fwe 85.53 91.27 83.07 84.33 77.13 80.33
qa_1 77.40 78.40 73.00 71.20 70.60 65.00
qa_2 60.00 64.20 56.80 61.40 49.40 59.60
avg 73.83 78.38 68.43 72.57 64.35 65.65
LLaMA vt 78.48 83.76 54.04 77.20 24.72 49.48
cwe 86.54 92.56 60.92 80.20 10.96 28.14
fwe 81.40 87.00 73.47 73.93 61.00 67.67
qa_1 76.60 74.40 74.20 74.40 68.00 67.80
qa_2 47.80 56.20 43.20 51.40 40.00 43.00
avg 74.16 78.78 61.17 71.43 40.94 51.22
Table 3: Performance comparison across sequence lengths.
A.4 Scaling Data to 3W
We collect 10,000 samples from each domain: books, arXiv, and code. We end up with 30,000 long document. The ratio of data with $K=2,4,6,8$ is $1:1:2:2$. We train on backbone model Qwen2.5-7B-Instruct-1M. As shown in Figure 9. The performance of our model can surpass baseline by about 10 points.
Figure 9: We report the performance of RULER when scaling the data to 30,000 samples on Qwen2.5-7B-Instruct-1M.

### v3

@@
 The improvements remain consistent across different backbone architectures, including Qwen2.5-7B-Instruct-1M and LLaMA-3.1-8B-Instruct, demonstrating that our method is not tied to a specific model family.
+Model #Data RULER LongBench
+Qwen2.5-7B-Instruct-1M
+   SFT 46K 64.5 33.2
+   Ours 14K 68.3 33.8
+LLaMA-3.1-8B-Instruct
+   SFT 46K 64.7 31.2
+   Ours 14K 61.2 30.4
+Table 1: Comparison between SFT and our reconstruction training on RULER (AVG) and LongBench v2 (Overall).
 Taken together, these results show that unsupervised document reconstruction via RLVR provides a scalable and effective mechanism for improving long-context capabilities.
@@
 Detailed scores on RULER can be found in Appendix A.3.
+Comparison to SFT
+Chen et al.
+(2026) employ a powerful teacher model to curate 46K context-specific QA pairs for supervised fine-tuning.
+They report results on LongBench v2 and only the QA subset of RULER.
+In Table 1, we compare our results with theirs.
+With only 14K training samples and without relying on context-specific QA pairs, our approach achieves superior performance on Qwen2.5-7B-Instruct-1M, while slightly lags behind on LLaMA-3.1-8B-Instruct.
 5.2 Dense vs. Sparse Reward
-Figure 4: Average scores of RULER. We compare performance of dense and sparse rewards.
 In our main experiments, we employ a dense reward that provides partial credit for partially correct reconstructions.
@@
 We hypothesize that sparse rewards are more likely to cause training instability due to the sparsity of positive rewards in the training process.
+Figure 4: Average scores on RULER. We compare the performance of dense and sparse rewards.
+Figure 5: Average scores on RULER. We compare the performance of different option length mixture ratios.
 5.3 Robustness to Option Mixture Ratios
-Figure 5: Average scores of RULER. We compare the performance of different option length mixture ratios.
 In our main experiments, we adopt the option mixture ratio for $K=2,4,6,$ and $8$ as $3:3:3:5$, which assigns only a moderate portion of training samples to small values of $K$ (e.g., KK=2).
@@
 5.4 Longer Documents Bring Consistent Improvement
-Figure 6: We report the average score of RULER
+Figure 6: We report the average score of RULER.
 During data curation of main experiment, we retain the longest 8,000, 3,000, and 3,000 documents from the book, arXiv, and code domains, respectively.
@@
 5.7 Analysis of Option Length
+In this part, we study the influence of sample difficulty on model performance.
+We randomly sample 5,000 documents from the training corpus and create four distinct training sets, where each set is constructed using a fixed option length $K\in\{2,4,6,8\}$, without mixing different values of $K$ within the same set.
+We train an independent model for each curated training set, where each model is associated with a specific $K$.
 Model Qwen LLaMA
@@
 Table 3: Average performance of different $K$ values on RULER.
-In this part, we study the influence of sample difficulty on model performance.
-Intuitively, a sample with more candidate options is more challenging, as the model must search over a larger permutation space and rely more heavily on global contextual cues.
-Specifically, we randomly down-sample 5,000 documents from the training corpus and create four distinct training sets, where each set is constructed using a fixed option length $K\in\{2,4,6,8\}$, without mixing different values of $K$ within the same set.
-We train an independent model for each curated training set, where each model is associated with a specific $K$.
 As shown in Table 3, we observe that model performance is relatively stable across different option lengths $K$.
@@
 Reinforcement Learning with Verifiable Rewards provides an objective and scalable framework for improving the reasoning capabilities of large language models by supervising them with ground-truth answers (Lambert et al., 2025; Guo et al., 2025).
-This paradigm has proven particularly effective in domains with unambiguous correctness criteria, including mathematics and programming.
 Previous work shows that RLVR can elevate models to expert-level reasoning performance by encouraging the discovery of correct internal reasoning trajectories through outcome-only rewards (OpenAI et al., 2024; Kim et al., 2025; Huang and Yang, 2025; Mo et al., 2025; Wang et al., 2026b).
@@
 In these settings, emergent behaviors, such as self-reflection, have been identified as important contributors to performance gains (Gandhi et al., 2025).
-Recent reasoning-oriented models, including OpenAI’s o1 series and DeepSeek, have popularized RLVR-based training pipelines and reinforcement learning algorithms such as GRPO.
 Long-Context Training of LLMs
@@
 Prior work pads questions with unrelated passages, shuffles document order, or fills contexts with irrelevant text to artificially increase sequence length (Trivedi et al., 2023).
-However, synthetic-data construction often introduces superficial complexity and fails to produce high-quality and challenging examples, limiting the effectiveness of the resulting long-context capabilities.
-Despite these limitations, improving long-context reasoning has become increasingly important due to the rapid emergence of agent applications (Zhao et al., 2024; Team et al., 2025a; b; Prabhakar et al., 2025; Gandhi et al., 2026).
-Thus, developing scalable and principled training methods that move beyond synthetic context expansion and provide effective supervision for long-context training is becoming urgent for advancing large language models.
+Improving long-context reasoning has become increasingly important due to the rapid emergence of agent applications (Zhao et al., 2024; Team et al., 2025a; b; Prabhakar et al., 2025; Gandhi et al., 2026).
 7 Conclusion
-In this work, we present an unsupervised reinforcement learning framework for improving the long-context capabilities of large language models through document reconstruction.
-By formulating reconstruction as a sequential decision-making problem with verifiable rewards derived directly from raw documents, our approach eliminates the need for manually curated long-context QA data or teacher-model supervision.
+In this work, we present an unsupervised reinforcement learning framework for improving the long-context capabilities of large language models.
+By formulating reconstruction as a sequential decision-making problem with verifiable rewards derived directly from raw documents, our approach eliminates the need for manually curated long-context data or teacher-model supervision.
 This enables a scalable and principled alternative to existing long-context training paradigms.
 Extensive experiments on RULER and LongBench v2 demonstrate that reconstruction-based RLVR effectively enhances long-context performance across multiple backbone models and context lengths.
-Beyond empirical gains, our findings highlight document structure itself as a valuable and underexplored supervision signal for long-context learning.
 We hope that this work motivates further research into unsupervised and self-supervised training objectives that leverage intrinsic structure in raw data and contributes to the development of more capable and scalable long-context language models.